OpenStudy (anonymous):

sean works for a heatin/cooling contractor. a customer wants to buy a humidifier for his living room. his room measures 15 ft. by 18 ft. his walls are 8 ft. tall. to determine how large a unit to buy, he must calculate the volume of the room. How many cubic feet of air are in the room? @smileyxl3

OpenStudy (smileyxl3):

Again, all you need is \[V = l \times w \times h\] so \[V = 15 \times 18 \times 8\]